OpenStudy (anonymous):

Determine which of the following binomials are factors of 6x^3-23x^2+29x-12 : 3x-2 3x+4 2x-3 3x+2 3x-4 If you could explain to me how you found the answers I would really appreciate it(;

OpenStudy (anonymous):

Substitute these 5 values of x into the polynomial f(x) to see if f(x) = 0. If so, then the binomial is a factor. Example 1. 3x - 2 = 0 -> x = 2/3. f(2/3) = -54 . Conclusion: (3x - 2) is not a factor. Continue with others.

OpenStudy (ranga):

Equate each of the answer choices to zero and find the root. 2/3, -4/3, etc.. Then do synthetic division of 6 - 23 + 29 - 12 by each of the roots. Those that give you a remainder of 0 are the factors.

OpenStudy (anonymous):

Thank you! :D